OpenStudy (anonymous):

Does anyone know the correct formula for linear thermal expansion for steel?

mathslover (mathslover):

Like what? Don't we already have a formula for linear thermal expansion of a solid as : \(L = L_0 (1 + \alpha \triangle T) \) Where L = Final Length of the solid. \(L_0\) = Initial Length of the Solid and \(\alpha\) = Coefficient of Linear Thermal Expansion. \(\triangle T\) = Change in Temperature.

OpenStudy (anonymous):

\[\color{green}{\Delta \text{L} = \alpha \text{L}_0 \Delta \text{T}}\]